OpenStudy (anonymous):

Two forces with magnitudes of 25 and 30 pounds act on an object at angles of 10° and 100° respectively. Find the direction and magnitude of the resultant force. Round to two decimal places in all intermediate steps and in your final answer. i need help figuring out how to do it, not just the answer please

OpenStudy (anonymous):

 we can add the x and y components separately and use our knowledge of whether x and y are positive or negative in the quadrant in which it lies. x components: 25*cos(10) - 30*cos(80) = Fx y components: 25*sin(10) + 30*sin(80) = Fy Magnitude [absolute value]: \[F = \left| \sqrt{Fx^{2} + Fy^{2}} \right|\] and the direction of F is the angle define by: \[\tan (\theta) = \frac{ Fy }{ Fx }\]